OpenStudy (anonymous):

2. Combine like terms 7? + 3? + 6? 4 − 4? 4 + 9? − 10

OpenStudy (anonymous):

can someone teach me how to combine like terms????

OpenStudy (anonymous):

brb im gonna do another question for now

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Combining like terms is the same as putting 2+2 together. for example: 2+2-3+4-3+2 = 6-6+4

OpenStudy (anonymous):

so is basically addition??

OpenStudy (anonymous):

its*

OpenStudy (anonymous):

In its own way, yes.

OpenStudy (anonymous):

ok so 16

OpenStudy (anonymous):

+ 13

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

The difference in combining like terms from addition is, rather then [2+2-4 = 0] like numbers or terms are pulled from anywhere in the equation so 2+2-4+2 would equal 6-4.

OpenStudy (anonymous):

ok im gonna watch a video on it but thanks

OpenStudy (anonymous):

Okay.